A slurry of flaked soybeans weighing a total of 100 kg contains 75 kg of inert solids and 25 kg of solutionwith 10 wt % oil and 90 wt % solvent hexane. This slurry is contacted with 100 kg of pure hexane ina single stageso that the value of N for the outlet underflow is 1.5 kg insoluble solid/kg solution retained. Calculate the composition of the underflow,L1 in weight percent oil.



1
Expert's answer
2021-03-30T07:49:56-0400

L0 = 25Kg

V2 = 100

yA0 = 0.1g

YA0 = 0

B = 75 kg (inert)

L0N2 = 25+100 = 125 kg = M

YA = Kg A/(KgA + Kg)

= 25×0.1/25×0.1 + 125(1-0.1)

= 0.1

L0yAO + V2 XAz = 25×0.1 + 100

= 125 Xm

Vm = 2.5/125 = 0.2

YA1 = 0.02

B= N0L0 = Nm = M = 3×25 = Nm × 125 = 75/125 = 0.6

B= N1L1 -> L1 = 75/1.5 = 50

= 50kg/h